OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x - 3y = -3 2x - 6y = -6 (2 points)Question 3 options: 1) -1 2) 0 3) 1 4) 2

OpenStudy (anonymous):

2x-6y=-6 2x=-6+6y x=-3+3y 5(-3+3y)-3y=-3 -15+15y-3y=-3 12y=12 y=1

OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + y = 3 2x + 2y = 10 (2 points)Question 5 options: 1) (1, 2) 2) (2, 3) 3) No Solutions 4) Infinitely Many Solutions